Solve for X, Y, and Z in the following systems of three equations using either substition or elimination methods:
a. X + 2Y + Z = 6
X + Y = 4
3X + Y + Z = 8
b. 10X + Y + Z = 12
8X + 2Y +Z = 11
20X - 10Y - 2Z = 8
c. 22X + 5Y + 7Z = 12
10X + 3Y + 2Z = 5
9X + 2Y + 12Z = 14

Answer by Alan3354(69443)

Solve for X, Y, and Z in the following systems of three equations using either substition or elimination methods:
X + 2Y + Z = 6
X + Y = 4
3X + Y + Z = 8
--------------
Eliminate one variable in 2 of the eqns by getting the same coefficients and eliminating.
Eliminate x
-----------
X + 2Y + Z = 6
X + Y = 4
--------------- Subtract
0x + y + z = 2 Eqn A
-----------------
X + Y = 4 times 3 --> 3x + 3y = 12
3X + Y + Z = 8
3x + 3y + 0z=12
----------------- Subtract
0x - 2y + z = -4 Eqn B
0x + y + z = 2 Eqn A
--------------- Subtract
-3y = -6
y = 2
------
z = 0
------
x = 2
=====================
Do the others the same way. email me for help, or to check your work @
gsihoutx@aol.com